OpenStudy (anonymous):

What is the GCF of the numerator and denominator in the following fraction? - 99/176 a) 3 b) 9 c) 11 d) They are relatively prime.

OpenStudy (anonymous):

99=3*3*11 176=2*2*2*2*11

OpenStudy (anonymous):

Thus, GCF=11

OpenStudy (anonymous):

got it?